OpenStudy (anonymous):

If \[ A=\begin{bmatrix} 1 & 2 & 0\\ 2 & -1& -2\\ 0 & -2& 1 \end{bmatrix} \] And it is given that \[P^{T}AP \] equals to a diagonal matrix, and that P is an orthogonal matrix, how can I find the matrix P?

OpenStudy (zarkon):

find the eigenspace

OpenStudy (anonymous):

but the eigenvalues include complex numbers.

OpenStudy (zarkon):

-3,1,3 are the eigenvalues

OpenStudy (zarkon):

http://alturl.com/c5aiw

OpenStudy (anonymous):

ohhh...thanks! i had a calculation mistake. thanks a lot!! :)

OpenStudy (zarkon):

np

OpenStudy (anonymous):

but I still have a problem... because the question wanted a transpose of P, after finding the eigenspace, to diagonalise the eigenvalue matrix, I need to get the inverse of the eigenspace matrix. If I get an inverse of it, then it isn't a transpose anymore.

OpenStudy (zarkon):

did you normalize the eigenvectors?

OpenStudy (anonymous):

but why do i have to normalise them?

OpenStudy (zarkon):

P is really just supposed to be a rotation matrix (an orthonormal matrix)

OpenStudy (anonymous):

wow... can I generalise it to say for all othonormal matrix that belongs to the eigenspace, the transpose is the same as the inverse?

OpenStudy (zarkon):

If you do not normalize then you will still diagonalize a but \[p^t\neq p^{-1}\]

OpenStudy (anonymous):

Can I generalise it to say all normalised othogonal matrix, its transpose is the same as its inverse?

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

it is really easy to prove since \[p^tp=I\] thus \[p^t=p^{-1}\]

OpenStudy (zarkon):

some take the definition of orthogonal matrix to be a square matrix A such that \[A^tA=I\] you should check your book to see how it defines it

OpenStudy (anonymous):

oh...thanks a lot for your help! thanks!

OpenStudy (zarkon):

np :)